Nominal Cost Help; Both questions, please.

Refer to table 1. Consider a basket of goods that includes 10 apples, 5 bananas, 3 kiwis, and 1 pineapple. At the prices given in the table, what is the nominal cost of the basket of goods in 2004? Table 1 ear apple baaa kiwi pineapple 2003 $4 $2 $5 $6 2004 $3 $2 $6 $8 Select one: O a. $80 o b. $66 O c. $73 O d. $40 Refer to table 1 again. Consider the same basket of goods (10 apples, 5 bananas, 3 kiwis, and 1 pineapple). At the prices given in the table, what is the rate of inflation between 2003 and 2004 (Use 2003 as the base year)? Table 1 apple banana kiwilpineapple 2003 $4 $2 $5 $6 2004 $3 $2 $6 $8 Select one: O a. 7% О с. 3% O d.-2%
Answer #1

Answer : 1) The answer is option b.

Nominal cost = current year price level * Basket of goods

For 2004 : Nominal cost = (3 * 10) + (2 * 5) + (6 * 3) + (8 * 1) = 30 + 10 + 18 + 8 = $66

Therefore, option b is correct.

2) The answer is option b.

CPI (Consumer price index) = [(Current year price * Basket of goods) / (Base year price * Basket of goods)] * 100

Here 2003 is the base year.

For 2003 : CPI = [{(4 * 10) + (2 * 5) + (5 * 3) + (6 * 1)} / {(4 * 10) + (2 * 5) + (5 * 3) + (6 * 1)}] * 100 = 100

For 2004 : CPI = [{(3 * 10) + (2 * 5) + (6 * 3) + (8 * 1)} / {(4 * 10) + (2 * 5) + (5 * 3) + (6 * 1)}] * 100 = [66 / 71] * 100 = 93

Inflation rate = [(CPI of 2004 - CPI of 2003) / CPI of 2003] * 100

=> Inflation rate = [(93 - 100) / 100] * 100

=> Inflation rate = - 7%

Therefore, option b is correct.